OpenStudy (anonymous):

5. A bag contains 6 silver and 5 gold coins. Which probability distribution represents the number of silver coins in a random draw of 4 coins?

OpenStudy (anonymous):

Does it say you put them back in or take them out and leave them out?

OpenStudy (uri):

@shabbir581 Aslamwalikum..:3

OpenStudy (anonymous):

probablity is 5...

OpenStudy (anonymous):

6/11×5/10×4/9×3/8

OpenStudy (uri):

NO.

OpenStudy (anonymous):

No who?

OpenStudy (anonymous):

X: no of silver coins in a randon draw of '4 ' coins so 'X can have values : 0, 1,2,3 OR 4 Accordingly, P(x) for each of these value will be P(x=0) = 5c4/ 11 C4 = 5/330 = 1/66 P(x=1) = 5c3*6C1/ 11 C4 =60/330 = 2/11 P(X=2) = 5c2*6C2/ 11 C4 = 150/330 = 5/11 P (X=3) = 5c1*6C3/ 11 C4 = 100/330 = 10/33 P (X= 4) = 5c0*6C4/ 11 C4 = 15/330 = 5/110 So Prob distribution is X: 0 1 2 3 4 P(X) = 5/330 60/330 150/330 100/330 15/330

OpenStudy (uri):

@shabbir581 Do you have answer choices..